OpenStudy (anonymous):

In a maths class of 26 students, each girl drew a triangle and each boy drew a rectangle. If there were 92 sides drawn by the students how many girls were in the class?

OpenStudy (anonymous):

3g + 4b = 92 g + b = 26

OpenStudy (anonymous):

u get it?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

first u have to know that a triangle has 3 sides g represent the number of girls in the class, each girl draws 3 sides, therefore all girls draw in total 3g sides

OpenStudy (anonymous):

same logic with 'b', which represents the number of boys

OpenStudy (anonymous):

do you get it now?

OpenStudy (anonymous):

3x + 4y = 92 ; x=no of sides of triangle used by girls and y= no of sides of rectangle used by boys. (as girls used three sides to form a triangle and boys used 4 sides to form a rectangle, adding both the quantities, we get the above equation :) ) x + y = 26; (addition of boys and girls) now solve them using equations in two variables